Vicente Wholesalers made sales using the following list prices and trade discounts. What amount will be recorded for each sale in the sales journal?

  1. List price of $800 and trade discount of 20 percent
  2. List price of $2,080 and trade discount of 40 percent
  3. List price of $1,500 and trade discount of 25 percent
Answer #1

Trade discount is a discount given by the seller to the buyer at the time of purchase of goods.

This is usually allowed by the sellers to attract more customers.

No record to be maintained for Trade discount.

It is normally dedcuted at the time of sale.

Journal entry is prepared based on net price

Following Journal entries will be prepared for above transactions

1.LP is $800,TD is 20%,Net price is 800-20%=640

Trade receivable .......Dr $640

Sales .................cr $640

2.LP is $2080,TD is 40%,Net priceis $ 2080-40%=1248

Trade receivable.......Dr $1248

Sales......................Cr $1248

3.LP is 1500,TD is 25%,Net price is $ 1500-25%=1125

Trade receivable......Dr $ 1125

Sales ....................Cr $ 1125